# Benchmarking Pretrained Molecular Embedding Models For Molecular Representation Learning

### v3

@@ Conclusions
CLAMP >ECFP 0.86 0.10 1.00 0.00
Atom Pair >ECFP 0.68 0.17 1.00 0.92
CDDD >ECFP 0.66 0.18 1.00 0.95
ChemBERTa >ECFP 0.60 0.19 0.96 1.00
ECFP >R-MAT 0.72 0.16 1.00 0.74
ECFP >MolBERT 0.79 0.13 1.00 0.19
ECFP >MoLFormer 0.80 0.12 1.00 0.11
ECFP >MAT 0.80 0.12 1.00 0.10
ECFP >ChemFM 0.89 0.07 1.00 0.00
ECFP >Mol2Vec 0.94 0.04 1.00 0.00
ECFP >TT 0.94 0.04 1.00 0.00
ECFP >GraphMVP 0.97 0.02 1.00 0.00
ECFP >COATI 0.98 0.01 1.00 0.00
ECFP >Chemformer 0.99 0.01 1.00 0.00
ECFP >MolR 0.99 0.01 1.00 0.00
ECFP >GIN 0.99 0.01 1.00 0.00
ECFP >Uni-Mol2 0.99 0.01 1.00 0.00
ECFP >ChemGPT 1.00 0.00 1.00 0.00
ECFP >GEM 1.00 0.00 1.00 0.00
ECFP >GROVER 1.00 0.00 1.00 0.00
ECFP >Uni-Mol 1.00 0.00 1.00 0.00
ECFP >SimSon 1.00 0.00 1.00 0.00
ECFP >SELFormer 1.00 0.00 1.00 0.00
ECFP >GraphFP 1.00 0.00 1.00 0.00
Table 5: Bayesian Bradley-Terry model statistics for the classification results.
Additional detailed results

### v4

@@ 14.2 Decision rules and ROPE
CLAMP >ECFP 0.62 0.10 1.00 0.84
R-MAT >ECFP 0.57 0.11 0.99 0.99
MolBERT >ECFP 0.52 0.11 0.73 1.00
CDDD >ECFP 0.50 0.11 0.55 1.00
ChemBERTa >ECFP 0.50 0.11 0.55 1.00
ECFP >MAT 0.51 0.11 0.64 1.00
ECFP >Atom Pair 0.55 0.10 0.94 1.00
ECFP >MoLFormer 0.60 0.10 1.00 0.93
ECFP >Mol2Vec 0.64 0.10 1.00 0.66
ECFP >TT 0.70 0.09 1.00 0.03
ECFP >ChemFM 0.74 0.08 1.00 0.00
ECFP >Uni-Mol2 0.76 0.08 1.00 0.00
ECFP >MolR 0.76 0.08 1.00 0.00
ECFP >COATI 0.77 0.07 1.00 0.00
ECFP >Chemformer 0.80 0.07 1.00 0.00
ECFP >GraphMVP 0.80 0.07 1.00 0.00
ECFP >Uni-Mol 0.86 0.05 1.00 0.00
ECFP >GIN 0.88 0.05 1.00 0.00
ECFP >GROVER 0.88 0.05 1.00 0.00
ECFP >GEM 0.92 0.03 1.00 0.00
ECFP >ChemGPT 0.92 0.03 1.00 0.00
ECFP >SimSon 0.92 0.03 1.00 0.00
ECFP >SELFormer 0.94 0.03 1.00 0.00
ECFP >GraphFP 0.99 0.01 1.00 0.00
Table 6: Bayesian Bradley-Terry model statistics for the classification results. ROPE = $[0.35,0.65]$
15 Additional detailed results